The alternator on a truck engine produces 350Wof electric power when it rotates at 4000 rev/min. If it is 95 percent efficient, what should the difference (in newtons) between the tensions in the tight and slack parts of its V-belt drive be? The diameter of the alternator’s pulley is 10 cm.


1
Expert's answer
2021-11-04T10:17:20-0400

1) Determine the torque:


"\tau=\frac P\omega=30/(4000/60)=0.45\text{ N}·\text{m}."

2) The tension applied to the pulley:


"T=\tau/r=0.45/0.05=9\text{ N}."

3) The difference is


"\Delta T=(1-\eta) T=(1-0.95)9=0.45\text{ N}."